# Wide-Field, High-Resolution Reconstruction in Computational Multi-Aperture Miniscope Using a Fourier Neural Network

### v1

@@ 3.3 SV-FourierNet achieves state-of-the-art reconstruction quality and speed
[…]
Figure 6: Comparison of SV-FourierNet with existing state-of-the-art networks on simulated and experimental datasets. (a) Reconstruction results from simulated cell and vasculature measurements. SV-FourierNet achieves the best PSNR and SSIM for both types of samples. (b) Reconstruction results from experimental measurements on 10 $\mu$m fluorescent beads and freely moving C. elegans. SV-FourierNet provides the best visual quality and the fastest inference speed. Two C. elegans are indicated by the white and yellow arrows in both the measurement and the reconstructed images. The numbers on the measurement index the views from each microlens. SV-FourierNet successfully reconstructs the C. elegans populations across the entire FOV, while other networks suffer from loss of information especially near the FOV edges (marked by the white arrows).
@@
To ensure a fair comparison, we employ the same loss function and training strategy across all models. Each model is trained on the same dataset for 48 hours. The models with the best SSIM are then chosen for the benchmark comparison. Moreover, the architecture of the enhancement network within MultiWienerNet is modified to be the same as that of SV-FourierNet. The detailed architecture of the benchmark networks can be found in Supplement 1.
[…]
MultiWienerNet improves reconstruction performance at the FOV edge through multiple initial Wiener deconvolutions. However, it fails to preserve high-resolution details, as shown in the zoomed-in areas. This underscores the advantage of learning inverse mapping of a spatially-varying model in the frequency domain over the spatial domain. The broad, smooth spectrum of the PSF, as opposed to the spatial domain’s sparse and sharp features, proves easier to learn and more resilient to potential errors possibly due to neural network’s spectral bias towards low-frequency features[36]. Frequency domain filters excel in addressing diverse frequency components, adeptly analyzing quasi-periodic signals from our MLA-based measurement and adapting to intricate noise characteristics. In contrast, Wiener deconvolution, optimized for uniform, additive white Gaussian noise, results in suboptimal initial reconstructions and the loss of fine details in complex scenarios.
[…]
@@
To showcase our system’s capability in imaging large-scale dynamic biological processes, we image colonies of freely moving C. elegans. Details about the sample preparation are provided in Supplement 1. The SV-FourierNet reconstruction [Fig. 7(a)] and concurrent widefield measurements [Fig. 7(b)] are compared.
@@ 3.5 Experimental demonstration on fixed brain section
To demonstrate our system’s capability for high-quality imaging of complex biological tissues, we image a 75 $\mu$m thick, weakly scattering brain section containing GFP-expressing neurons in the bed nucleus of the stria terminalis (BNST). The BNST, crucial in regulating emotional and stress responses, is bilaterally located on either side of the brain, approximately 2.5 mm apart, with each region extending about 2 mm. These regions are characterized by their intricate spatial structure and varying neuron density levels (detailed sample information is in Supplement 1).
First, we evaluate the full-FOV SV-FourierNet reconstruction by comparing it with a 2$\times$ widefield measurement. Upon visual inspection, SV-FourierNet reliably reconstructs the bilateral structure of the labeled brain regions, as shown in Fig. 8(a). We then compare the reconstructions of four zoomed-in regions from both sides of the brain section with 10$\times$ widefield measurements (shown in Fig. 8(b)). SV-FourierNet accurately reconstructs neuronal structures from both densely and sparsely populated areas. These results demonstrate that SV-FourierNet, trained entirely on simulated data, can generalize to complex brain tissues, provide cellular-resolutions across the entire brain section, and is robust against weak tissue scattering.
@@ 4 Conclusion
We present a computational multi-aperture miniscope augmented with SV-FourierNet that performs single-shot high-resolution wide-FOV imaging. Our system is simple and compact, using only a single MLA for imaging. Unlike studies focusing on exploiting the 3D imaging capability with an MLA[9], we introduce a novel computational framework to significantly expand the FOV while ensuring uniform high resolution throughout. Our main innovation is SV-FourierNet, which employs multiple learnable Fourier filters to synthesize information over an extended FOV, addressing spatially-varying aberrations and view multiplexing inherent to our system. Furthermore, we demonstrate a novel utility of the network’s ERF in elucidating the mapping function it learns. We show that SV-FourierNet learns a physically meaningful reconstruction function for the spatially-varying imaging model, achieving uniform resolution throughout the extended FOV. Trained exclusively on simulated data, SV-FourierNet exhibits robust generalization to diverse experimental samples, as demonstrated on resolution targets, fluorescent beads, live C. elegans, and brain tissue sections. This underscores SV-FourierNet’s potential for a wide array of applications within biomedical research and beyond.
[…]
For the first time to our knowledge, we establish a connection between the ERF of a deconvolution network and the underlying physical imaging model. This enables a straightforward visualization of the function learned by the network, alleviating the “black box” aspect often associated with the deep learning networks. Furthermore, the ERF can offer a promising tool to inform network design, allow for predictions of network performance and alignment between the trained deconvolution model and the physical optical system by visualizing deviations of the ERFs from the physical PSFs.
@@
Acknowledgments The authors thank Boston University Shared Computing Cluster for proving the computational resources, Ian G. Davison, Kevin J Monk, and Brett T DiBenedictis for providing the brain slices, as well as Christopher A Gabel and Andrew S. Chang for providing the C. elegans sample. We also thank Boston University Computational Imaging Systems Lab group members for insightful discussions and Adelina Chau and Jamin Xie for their assistance with benchmark network tuning.

### v2

@@ 3.3 SV-FourierNet achieves state-of-the-art reconstruction quality and speed
To demonstrate the superior performance of SV-FourierNet, we compare the reconstructions from our network with the LSV-model-based ADMM algorithm and the existing state-of-the-art reconstruction networks, including CM2Net[14] and MultiWienerNet[21]. We demonstrate that SV-FourierNet achieves the best reconstruction quality with the fastest inference speed on diverse types of samples in simulation and experiments.
Figure 6: Comparison of SV-FourierNet with existing state-of-the-art networks and LSV model-based algorithm (ADMM) on simulated and experimental datasets. (a) Reconstruction results from simulated cell measurement. SV-FourierNet achieves the best PSNR and SSIM. (b) Reconstruction result from experimental measurement on a resolution target. SV-FourierNet provides the best spatial resolution (highest spatial resolution achieved by each method is marked by the yellow box) and the fastest inference speed.
@@
To ensure a fair comparison between the deep learning models, we employ the same loss function and training strategy across all models. Each model is first trained on the same dataset for 48 hours. The optimal models achieving the highest SSIM are then chosen for the benchmark comparison (see details in Section 7 in Supplement 1). Moreover, the architecture of the enhancement network within MultiWienerNet is modified to be the same as that of SV-FourierNet. The detailed architecture of the benchmark networks can be found in Section 7 in Supplement 1.
We compare the reconstruction results from CM2Net, MultiWienerNet, SV-FourierNet and LSV ADMM algorithm on both simulated testing data and experimental measurements, shown in Fig. 6. From metrics evaluation, SV-FourierNet achieves the best PSNR and image quality (additional quantitative analysis on simulated testing data is provided in Section 7 in Supplement 1). From the resolution target reconstruction results, SV-FourierNet demonstrates superior resolution compared to CM2Net and MultiWienerNet and is comparable to the LSV ADMM algorithm. Notably, at the edge of the FOV, SV-FourierNet maintains consistently high resolution, whereas other techniques exhibit degraded resolution, as detailed in Section 8A of Supplement 1. By visual inspection, CM2Net suffers from severe ghosting artifacts and a degradation in performance at the FOV’s peripheral regions. This is attributed to CM2Net’s patch-based training strategy, which neglects the strong shift variance between patches. Similar artifacts are observed in reconstructions from model-based deconvolution algorithm using a spatially invariant model (see details in Section 2A in Supplement 1), demonstrating the importance of accounting for spatially-varying degradation in the reconstruction process.
MultiWienerNet improves reconstruction performance at the FOV edge through multiple initial Wiener deconvolutions. However, it fails to preserve high-resolution details and suffers from pronounced background artifacts compared to SV-FourierNet. In addition to superior image quality, SV-FourierNet outperforms other methods with its enhanced reconstruction speed and less computational load. By leveraging the fast Fourier transform’s computational efficiency, SV-FourierNet achieves a 20$\times$ speedup compared to CM2Net. Furthermore, by utilizing the Fourier transform’s symmetry property for real-valued signals and avoiding the additional computational costs associated with Wiener deconvolution, SV-FourierNet reduces 2$\times$ memory consumption and achieves a 2$\times$ speedup over MultiWienerNet (See additional comparisons for experimental and simulation results in Section 7 in Supplement 1).
To understand the superior performance of SV-FourierNet and our training strategy, we conducted extensive ablation studies, leading to two main insights: the impact of PSF/filter initialization and the choice of learning domain (spatial vs Fourier). First, initializing with physical PSFs, regardless of the learning domain, results in reduced resolution and pronounced noisy artifacts (see Section 6D in Supplement 1). This is because physical PSF initialization introduces unwanted locality biases and inherent noise in the PSF measurements. Inspecting the learned filters in MultiWienerNet initialized with physical PSFs (Section 6C in Supplement 1) reveals that the learning process primarily adds non-interpretable noisy features to the filters, similar to those reported in [21]. This argument is further corroborated by the ERF of MultiWienerNet (Section 5B in Supplement 1), which shows severe noisy artifacts in the learned functions. Similar artifacts are observed when SV-FourierNet is trained with the same physical PSF initialization, instead of the random initialization used in our method. Replacing the physical PSF initialization with the same random initialization allows MultiWienerNet to converge to more physically meaningful filters, similar to those learned by SV-FourierNet, but with reduced resolution. These results lead to our second insight: learning in the spatial domain generally results in inferior resolution compared to learning in the Fourier domain. This can be explained by two factors: the smooth PSF spectrum improves the conditioning of the learning problem, and Fourier domain learning eliminates an unnecessary Fourier transform operation in the spatial domain approach, reducing noise and error propagation.
@@
To showcase our system’s capability in imaging large-scale dynamic biological processes, we image colonies of freely moving C. elegans. Details about the sample preparation are provided in Section 9D in Supplement 1. The SV-FourierNet reconstruction [Fig. 7(a)] and concurrent widefield measurements [Fig. 7(b)] are compared.
@@ 3.5 Experimental demonstration on fixed brain section
To demonstrate our system’s capability for high-quality imaging of complex biological tissues, we image a 75 $\mu$m thick, weakly scattering brain section containing GFP-expressing neurons in the bed nucleus of the stria terminalis (BNST). The BNST, crucial in regulating emotional and stress responses, is bilaterally located on either side of the brain, approximately 2.5 mm apart, with each region extending about 2 mm. These regions are characterized by their intricate spatial structure and varying neuron density levels (detailed sample information is in Section 9C in Supplement 1).
First, we evaluate the full-FOV SV-FourierNet reconstruction by comparing it with a 2$\times$ widefield measurement. Upon visual inspection, SV-FourierNet reliably reconstructs the bilateral structure of the labeled brain regions, as shown in Fig. 8(a). We then compare the reconstructions of four zoomed-in regions from both sides of the brain section with 10$\times$ widefield measurements (shown in Fig. 8(b)). SV-FourierNet accurately reconstructs neuronal structures from both densely and sparsely populated areas (An additional overlay comparison between SV-FourierNet reconstruction and 2$\times$ widefield measurement is provided in Section 9E of Supplement 1). These results demonstrate that SV-FourierNet, trained entirely on simulated data, can generalize to complex brain tissues, provide cellular-resolutions across the entire brain section, and is robust against weak tissue scattering.
@@ 4 Conclusion
We present a computational multi-aperture miniscope augmented with SV-FourierNet that performs single-shot high-resolution wide-FOV imaging. Our system is simple and compact, using only a single MLA for imaging. Unlike studies focusing on exploiting the 3D imaging capability with an MLA[13], we introduce a novel computational framework to expand the FOV while ensuring uniform high resolution throughout. Our main innovation is SV-FourierNet, which employs multiple learnable Fourier filters to synthesize information over an extended FOV, addressing spatially-varying aberrations and view multiplexing inherent to our system. Furthermore, we demonstrate a novel utility of the network’s ERF in elucidating the mapping function it learns. We show that SV-FourierNet learns a physically meaningful reconstruction function for the spatially-varying imaging model, achieving uniform resolution throughout the extended FOV. Trained exclusively on simulated data, SV-FourierNet exhibits robust generalization to diverse experimental samples, as demonstrated on resolution targets, fluorescent beads, live C. elegans, and brain tissue sections. This underscores SV-FourierNet’s potential for a wide array of applications within biomedical research and beyond.
We show SV-FourierNet’s superior capability to demultiplex cross-talks between regular MLAs and tackle severe spatially-varying degradation in 2D imaging scenarios. This principle also holds promise for expanding the FOV in 3D imaging applications. For example, the Fourier lightfield microscope employs a field stop to prevent view multiplexing between lenses, which inevitably reduces the FOV in proportion to the lens count, thus limiting its imaging capacity for large-scale biological processes[40]. Our strategy suggests the possibility of removing the field stop and harnessing view multiplexing to increase the FOV without compromising spatial resolution. In 3D multi-aperture microscopes with view multiplexing, like CM2, existing 3D reconstruction algorithms suffer from resolution degradation near the FOV edges[14]. Extending SV-FourierNet to 3D imaging may substantially improve the imaging volume with uniform 3D resolution. Thus, we anticipate that SV-FourierNet could offer a novel paradigm for single-shot wide-FOV high-resolution volumetric imaging.
SV-FourierNet effectively learns global shift-variant inverse filters in the frequency domain, which not only significantly enhances image quality with higher spatial resolution but also substantially reduces the computational burden compared to spatial domain learning. This reduction is crucial for model scalability, especially when handling large volumes of data. To further decrease model size and computational load, a promising direction involves exploring the sparse nature of the system’s PSFs. Such reduction in the model size will enable SV-FourierNet to be adapted to more challenging applications, such as volumetric shift-variant deconvolution.
For the first time to our knowledge, we establish a connection between the ERF of a deconvolution network and the underlying physical imaging model. This enables a straightforward visualization of the function learned by the network, alleviating the “black box” aspect often associated with the deep learning networks. Furthermore, the ERF can offer a promising tool to inform network design, allow for assessing network performance and alignment between the trained deconvolution model and the physical optical system by visualizing deviations of the ERFs from the physical PSFs.
@@
Acknowledgments The authors thank Boston University Shared Computing Cluster for proving the computational resources, Ian G. Davison, Kevin J Monk, and Brett T DiBenedictis for providing the brain slices, as well as Christopher A Gabel and Andrew S. Chang for providing the C. elegans sample. We also thank BU CISL group members for insightful discussions and Adelina Chau and Jamin Xie for their assistance with benchmark network tuning.